8) The city of Hope has a labor fonce of 1000. Twenty people lose their jobs each month and remain8)_ unemployed for exactly one month before finding jobs. On January 1, May 1, and September 1 of each year, 50 people lose their jobs for a period of four months before finding new jobs. What is the average duration of an unemployment spell? A) 3.43 months B) 2.85 months C) 3.14 months D) 1.25 months E) 2.15 monthsshow WORK!
Answer #1

Answer

Unemployment rate in any given month   70/1000 = 7%

Unemployment spells =

Short spells: 20 each month × 12 months = 240.

Long spells: 50 each × 3 = 150.

Total spells = 240 + 150 = 390. There are more number of short spells

   Total Duration of all spells = (240 spells × 1 (month)) 240 (months) + (150 × 4 (months)) 600 months = 840 months.

AD = TD/Total spells AD = Average duration    TD = Total duration

= 840/390 = 2.15 months.